# The Unbounded Denominators Conjecture

### v3

@@ 2.5.4. A potential-theoretic generalization
The path with §§2.1,2.5 leads straightforwardly to an extension in potential theory, which we formulate without detailing a proof. Consider a compact subset $K\subset\overline{D(0,1)}$ with transfinite diameter $d(K):=\mathrm{cap}(K,[\infty])$. The equilibrium measure $\mu_{K}$ is the unique probability measure supported by $K$ that attains the lowest possible logarithmic energy of
$\iint_{K\times K}\log{\frac{1}{|x-y|}}\,\mu_{K}(x)\,\mu_{K}(y)=-\log{d(K)}.$
@@
The result is still more general than 2.5.4, and the restriction here to $\mathbf{Z}\llbracket t\rrbracket$ expansions was chosen as minimal for our application to noncongruence modular forms. In a sequel work we will generalize our integrated holonomy rank bound, in particular to the case of $\mathbf{Q}\llbracket t\rrbracket$ formal functions, and study its applications to transcendence theory. With regard to the latter, it is of some interest to inquire about the optimal numerical constant that could take the place of the coefficient $e$ in (2.5.24).
@@ 3. Our approach to the Unbounded Denominators Conjecture
In this section, we lay out our main approach to the unbounded denominators conjecture. This will reduce the proof to a number of independent results in group theory, complex geometry, and complex analysis which we take up in sections §4, §5, and §6. Our main idea is to use our arithmetic holonomicity theorems to prove the following:
@@
Let $t:=q^{1/N}=e^{\pi i\tau/N}$. We use Theorem 2.0.5 with $U:=\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$, $p(x):=x^{N}$ and
@@
The integrality and cusp widths conditions in the definition of the $\mathbf{Q}(\lambda)$-vector space $R_{2N}$ entail a basis of $R_{2N}$ made of elements of the ring $\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$. More precisely, for a modular function $f$ with Fourier expansion at $i\infty$ lying in $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, by our choice of $t=q^{1/N}$ and $x(t)=(\lambda(q)/16)^{1/N}$ we have $x^{*}f\in\mathbf{Z}\llbracket t\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, on defining $x^{*}f$ as the formal $x$-expansion of $f(q)=f(q(x))\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}\subset\mathbf{Q}\llbracket x\rrbracket$. As $f$ is a regular function on some affine modular curve $Y$ over $\overline{\mathbf{Q}}$ which admits a finite étale map to $Y(2)_{\overline{\mathbf{Q}}}$, there exists a (minimal) nonzero linear differential operator $L$ over $\overline{\mathbf{Q}}(\lambda)$ such that $L(f)=0$ and all singularities $\lambda\neq 0,1,\infty$ of $L$ in $\mathbf{C}\smallsetminus\{0,1\}$ have trivial local monodromy. Moreover, our assumption on the cusp widths dividing $2N$ implies that a local coordinate in a small neighborhood of each cusp of $Y$ above $\lambda=0$ can be chosen to be the lift of some (positive integer) power of $x=(\lambda/16)^{1/N}$. This means that the pullback of $L$ to $U\smallsetminus\{0\}=\mathbf{C}^{\times}\smallsetminus 16^{-1/N}\mu_{N}$ admits a full set of meromorphic solutions in some sufficiently small neighborhood of $x=0$, i.e. has a trivial local monodromy around $x=0$. Therefore $f\in\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$, and $R_{2N}\subset\mathcal{V}(U,x(t),\mathbf{Z})$.
@@
By (1) and the choice of radius $r=1-AN^{-3}/2$, we have
@@
Corollary 2.0.5 now upper-bounds our requisite dimension as
$\dim_{\mathbf{Q}(x^{N})}\mathcal{V}(U,x(t),\mathbf{Z})\leq e\cdot\frac{\int_{|z|=1-A/(2N^{3})}\log^{+}{|F_{N}^{N}|}\,\mu_{\mathrm{Haar}}}{cN^{-3}}.$
The requisite bound by $O(N^{3}\log{N})$ now results follows from (2) with the choice $B:=A/2$. ∎
@@
We may also prove this proposition by using Theorem 2.0.1 directly. Using the notation in the proof, let $Y^{\prime}$ denote the modular curve $Y$ with all the cusps above $0\in Y(2)\cup\{0\}$ filled in. The fiber product $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ with its natural map to $U$ is a covering map (one can check this claim locally; the assumption on cusp widths is used to prove that $0\in U$ is not ramified). Therefore, the universal covering map $D(0,1)\rightarrow U$ factors through $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ and thus we obtain a map $D(0,1)\rightarrow Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U\rightarrow Y^{\prime}$ (the second map is the natural map) such that its composition with $Y^{\prime}\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$ is the map $D(0,1)\rightarrow U\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$. Thus $f\circ\varphi$ is also given by the natural pullback of $f$ from $Y^{\prime}$ to $D(0,1)$ and thus it is holomorphic over $D(0,1)$ as far as $f$ is holomorphic at all cusps in $Y^{\prime}$, which can be achieved by multiplying $f$ with a suitable power of $\lambda$. Thus we verify the analyticity property in Theorem 2.0.1. The rest of the proof is the same as the proof above.
@@ 3.1. A guide to the proof of the main theorem
    In §6 we solve this uniformity problem on the abstract grounds of Nevanlinna theory. We combine the crude growth bound on $|F_{N}|$ with a version of Nevanlinna’s lemma on the logarithmic derivative to prove our requisite uniform upper estimate on the mean proximity function $m(r,F_{N})=\int_{|z|=r}\log^{+}{|F_{N}|}\,\mu_{\mathrm{Haar}}$.
@@ 4.1. Wohlfahrt Level
We begin by recalling a notion of level for noncongruence subgroups due to Wohlfahrt [Woh64]. Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup. (Many of the arguments of this section do not require this hypotheses but since it is satisfied for our applications we assume it to avoid unnecessary distractions.) The group consisting of the two matrices $\pm I$, where $I=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right)$, will be denoted by $E$. The group $\mathrm{SL}_{2}(\mathbf{Z})$ acts via Möbius transformations both on the upper half plane $\mathbf{H}$ and the extended upper half plane $\mathbf{H}^{*}=\mathbf{H}\cup\mathbf{P}^{1}(\mathbf{Q})$. The action of $\mathrm{SL}_{2}(\mathbf{Z})$ on $\mathbf{P}^{1}(\mathbf{Q})$ is transitive. It follows that if a non-trivial element $\gamma\in G$ fixes an element $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$, then $\gamma$ has the form $\pm MU^{m}M^{-1}$, where $M\in\mathrm{SL}_{2}(\mathbf{Z})$, $M\infty=\zeta$, and
@@
Let $G$ and $H$ be finite index subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$ both containing $E$. Suppose that $L(G)$ and $L(H)$ both divide $N$, then any cusp of $G\cap H$ also has cusp width dividing $N$.
@@
    $\widetilde{\gamma}=(A^{-1}BA)\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)(A^{-1}BA)^{-1}=C\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)C^{-1},$
@@
    and thus in $\widetilde{H}$. Hence either $\widetilde{h}=\widetilde{\gamma}$ if $\widetilde{\gamma}$ lies in $\mathrm{SL}_{2}(\mathbf{Z})$ or $\widetilde{h}=(\widetilde{\gamma})^{p}$. In particular, the cusp width at this cusp is either $n$ or $n/p$ and certainly divides $N$ and hence also $Np$.
@@ 4.2. Modular Forms
For an even integer $N$, we will consider the following spaces of modular functions with rational coefficients generated by forms with bounded denominators, that is, subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}=\mathbf{Q}\llbracket q^{1/N}\rrbracket[1/q]$ generated by elements of $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes\mathbf{Q}$ as a $\mathbf{Q}(\lambda)$-vector space.
@@
    Let $M_{N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions on the modular curve $Y(N)=\mathbf{H}/\langle E,\Gamma(N)\rangle$ with coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$.
@@
    Let $R_{N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions with coefficients in $\mathbf{Q}$, bounded denominators at the cusp $\zeta=i\infty$, and cusp widths dividing $N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$.
(The vector space $R_{N}$ was also defined in Proposition 3.0.1 but we repeat the definition here for convenience.) For example, the (weight 0) holomorphic modular forms on $Y(2)$ are given by $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$, and the $\mathbf{Q}(\lambda)$-vector space generated by such elements inside $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$ is $M_{2}=\mathbf{Q}(\lambda)$.
@@
There is a containment $M_{N}\subset R_{N}$, and $M_{N}$ and $R_{N}$ have finite dimensions over $M_{2}=\mathbf{Q}(\lambda)$.
@@
Let $f$ be a holomorphic modular function on $Y(N)$, that is, a meromorphic function on the compact modular curve $X(N)$ whose poles are all at the cusps. Assume also that $f$ has coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$, Then the modular form $f\Delta(\tau)^{m}$ is holomorphic at the cusps for sufficiently large $m$. Moreover, $f\Delta(\tau)^{m}$ has coefficients in $\mathbf{Q}$. It follows from [Shi71, Theorem 3.52] that $f\Delta(\tau)^{m}$ has bounded denominators, and this implies the same for $f$, and thus there is a containment $M_{N}\subset R_{N}$.
The second claim follows from Corollary 2.0.5 and the remark (cf. the second paragraph of §1.1.3) that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$. Indeed, $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is a well-defined holomorphic map with unit derivative at the origin, and hence by Schwarz’s lemma the universal covering $D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has derivative strictly larger than $1$ in absolute value. (Later, in Theorem 5.1.4 below, we will exactly compute this latter derivative.) ∎
@@
The vector spaces $M_{N}$ and $R_{N}$ are fields. The space $M_{N}$ may be identified with the field of rational functions on the modular curve $Y(N)/\mathbf{Q}$. There are injective algebra maps
$M_{2}\rightarrow M_{N}\rightarrow R_{N}.$
The space $R_{N}$ is invariant under a normal finite index subgroup $G_{N}\subset\langle E,\Gamma(N)\rangle\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(G_{N})=N$.
@@
Note that $M_{N}$ and $R_{N}$ are subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$, which is a domain. Hence if $M_{N}$ and $R_{N}$ are rings then they are also integral domains, and any integral domain which has finite dimension over a field is itself a field.
The curve $Y(N)$ has a standard model over $\mathbf{Q}$ (as a moduli space of elliptic curves $C$ with a given symplectic isomorphism $C[N]\simeq\mathbf{Z}/N\mathbf{Z}\oplus\mu_{N}$) such that the cusp $\zeta=i\infty$ is defined over $\mathbf{Q}$, and the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$ on the global sections of $Y(N)$ is compatible with the $q$-expansion map. It follows that the set of generators of $M_{N}$ is closed under addition and multiplication and hence that $M_{N}$ is a ring, and thus a field. Moreover, $M_{N}$ contains the global sections of the (affine) curve $Y(N)/\mathbf{Q}$, and hence $M_{N}$ must be the function field of $Y(N)/\mathbf{Q}$.
The vector space $R_{N}$ is generated by holomorphic modular forms with bounded denominators at $\zeta=i\infty$. To show $R_{N}$ is a ring, it suffices to show that the product of any two such generators $g$ and $h$ is also a generator. Certainly $gh$ is a holomorphic modular form with rational coefficients and bounded denominators, so it suffices to show that the cusp width still divides $N$. But we may assume that $g$ and $h$ are invariant under finite index subgroup $G,H\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$, and thus $gh$ is invariant under $G\cap H$. It follows from Lemma 4.1.3 that $L(G\cap H)$ also has Wohlfahrt level dividing $N$.
Since $R_{N}$ is finite over $M_{2}$, it is generated by a finite number of basis elements each of which is invariant under some finite index subgroup $\Phi\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(\Phi)$ dividing $N$. The intersection of all these groups still has finite index and level $N$ by Lemma 4.1.3, and then we take $G_{N}$ to be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in this intersection, which also has $L(G_{N})=N$ by Lemma 4.1.4. ∎

### v4

@@ 2.5.27. A potential-theoretic generalization
The path with §§ 2.1 and 2.5 leads straightforwardly to an extension in potential theory, which we formulate without detailing a proof. Consider a compact subset $K\subset\overline{D(0,1)}$ with transfinite diameter $d(K)$ and equilibrium measure $\mu_{K}$. This means that the logarithmic energy functional satisfies
$\iint_{K\times K}\log{\frac{1}{|x-y|}}\,\mu(x)\,\mu(y)\geq-\log{d(K)}$
for all probability measures $\mu$ supported by $K$, and the equality is attained if and only if $\mu=\mu_{K}$. See, for example, [Kir05] for these definitions and their basic properties, including the relation to capacitance.
@@
The result is still more general than 2.5.27, and the restriction here to $\mathbf{Z}\llbracket t\rrbracket$ expansions was chosen as minimal for our application to noncongruence modular forms. In a sequel work we will generalize our integrated holonomy rank bound, in particular to the case of $\mathbf{Q}\llbracket t\rrbracket$ formal functions, and study its applications to transcendence theory. With regard to the latter, it is of some interest to inquire about the optimal numerical constant that could take the place of the coefficient $e$ in (2.5.28).
@@ 3. Our approach to the Unbounded Denominators Conjecture
In this section, we lay out our main approach to the unbounded denominators conjecture. This will reduce the proof to a number of independent results in group theory, complex geometry, and complex analysis which we take up in §§ 4, 5, and 6. Our main idea is to use our arithmetic holonomicity theorems to prove the following:
@@
Let $t:=q^{1/N}=e^{\pi i\tau/N}$. We use Corollary 2.0.5 with $U:=\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$, $p(x):=x^{N}$ and
@@
The integrality and cusp widths conditions in the definition of the $\mathbf{Q}(\lambda)$-vector space $R_{2N}$ entail a basis of $R_{2N}$ made of elements of the ring $\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$. More precisely, for a modular function $f$ with Fourier expansion at $i\infty$ lying in $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, by our choice of $t=q^{1/N}$ and $x(t)=(\lambda(q)/16)^{1/N}$ we have $x^{*}f\in\mathbf{Z}\llbracket t\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, on defining $x^{*}f$ as the formal $x$-expansion of $f(q)=f(q(x))\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}\subset\mathbf{Q}\llbracket x\rrbracket$. As $f$ is a regular function on some affine modular curve $Y$ over $\overline{\mathbf{Q}}$ which admits a Galois finite étale map to $Y(2)_{\overline{\mathbf{Q}}}$, a minimal-order nonzero linear differential operator $L$ over $\overline{\mathbf{Q}}(\lambda)$ with $L(f)=0$ has trivial local monodromies around any $\lambda\neq 0,1,\infty$. Indeed, by the minimality of $L$, this amounts to analytically continuing the algebraic function $f(\lambda)\in\mathbf{C}\llbracket\lambda^{1/N}\rrbracket$ along all paths in $Y(2)_{\mathbf{C}}=\mathbf{C}\smallsetminus\{0,1\}$; this is for instance since, by the lifting property for covering maps, every holomorphic map $D(0,1)\to Y(2)_{\mathbf{C}}=\mathrm{Spec}\,\mathbf{C}[\lambda,1/\lambda,1/(1-\lambda)]$ based at $0\mapsto y_{0}\in Y(2)_{\mathbf{C}}$ lifts to a holomorphic map $D(0,1)\to Y_{\mathbf{C}}$ based at an arbitrary fiber point of $y_{0}$ under the covering $Y\to Y(2)$. Moreover, our assumption on the cusp widths dividing $2N$ implies that a local coordinate in a small neighborhood of each cusp of $Y$ above $\lambda=0$ can be chosen to be the lift of some (positive integer) power of $x=(\lambda/16)^{1/N}$. This means that the pullback of $L$ to $U\smallsetminus\{0\}=\mathbf{C}^{\times}\smallsetminus 16^{-1/N}\mu_{N}$ admits a full set of meromorphic solutions in some sufficiently small neighborhood of $x=0$, i.e. has a trivial local monodromy around $x=0$. Therefore $f\in\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$, and $R_{2N}\subset\mathcal{V}(U,x(t),\mathbf{Z})$.
@@
By assumption 3.0.1(1) of Proposition 3.0.1 and the choice of radius $r=1-AN^{-3}/2$, we have
@@
Corollary 2.0.5 now gives the upper bound
(3.0.5) $\dim_{\mathbf{Q}(x^{N})}\mathcal{V}(U,x(t),\mathbf{Z})\leq e\cdot\frac{\int_{|z|=1-A/(2N^{3})}\log^{+}{|F_{N}^{N}|}\,\mu_{\mathrm{Haar}}}{cN^{-3}}.$
From assumption (2) of Proposition 3.0.1 (with the choice $B:=A/2$) together with the identity $\log^{+}|F_{N}^{N}|=N\log^{+}|F_{N}|$, we have
$e\cdot\frac{\int_{|z|=1-A/(2N^{3})}\log^{+}{|F_{N}^{N}|}\,\mu_{\mathrm{Haar}}}{cN^{-3}}\ll_{B}e\cdot\frac{\displaystyle{N\cdot\frac{\log{N}}{N}}}{cN^{-3}}=O(N^{3}\log N),$
which, combined with equation (3.0.5), is the desired upper bound. ∎
@@
We may also prove this proposition by using Theorem 2.0.1 directly. Using the notation in the proof of Proposition 3.0.1, let $Y^{\prime}$ denote the modular curve $Y$ with all the cusps above $0\in Y(2)\cup\{0\}$ filled in. The fiber product $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ with its natural map to $U$ is a covering map (one can check this claim locally; the assumption on cusp widths is used to prove that $0\in U$ is not ramified). Therefore, the universal covering map $D(0,1)\rightarrow U$ factors through $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U$ and thus we obtain a map $D(0,1)\rightarrow Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}}U\rightarrow Y^{\prime}$ (the second map is the natural map) such that its composition with $Y^{\prime}\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$ is the map $D(0,1)\rightarrow U\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$. Thus $f\circ\varphi$ is also given by the natural pullback of $f$ from $Y^{\prime}$ to $D(0,1)$ and thus it is holomorphic over $D(0,1)$ as far as $f$ is holomorphic at all cusps in $Y^{\prime}$, which can be achieved by multiplying $f$ with a suitable power of $\lambda$. Thus we verify the analyticity property in Theorem 2.0.1. The rest of the proof is the same as above.
@@ 3.1. A guide to the proof of the main theorem
    In § 6, we solve this uniformity problem on the abstract grounds of Nevanlinna theory. We combine the crude growth bound on $|F_{N}|$ with a version of Nevanlinna’s lemma on the logarithmic derivative to prove our requisite uniform upper estimate on the mean proximity function $m(r,F_{N})=\int_{|z|=r}\log^{+}{|F_{N}|}\,\mu_{\mathrm{Haar}}$.
@@ 4.1. Wohlfahrt Level
We begin by recalling a notion of level for noncongruence subgroups due to Wohlfahrt [Woh64]. Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup. (Many of the arguments of this section do not require this hypotheses but since it is satisfied for our applications we assume it to avoid unnecessary distractions.) The group consisting of the two matrices $\pm I$, where $I=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right)$, will be denoted by $E$. The group $\mathrm{SL}_{2}(\mathbf{Z})$ acts via Möbius transformations both on the upper half plane $\mathbf{H}$ and the extended upper half plane $\mathbf{H}^{*}=\mathbf{H}\cup\mathbf{P}^{1}(\mathbf{Q})$. The action of $\mathrm{SL}_{2}(\mathbf{Z})$ on $\mathbf{P}^{1}(\mathbf{Q})$ is transitive. It follows that if a nontrivial element $\gamma\in G$ fixes an element $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$, then $\gamma$ has the form $\pm MU^{m}M^{-1}$, where $M\in\mathrm{SL}_{2}(\mathbf{Z})$, $M\infty=\zeta$, and
@@
Let $G$ and $H$ be finite index subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$ both containing $E$. Suppose that $L(G)$ and $L(H)$ both divide $N$. Then any cusp of $G\cap H$ also has cusp width dividing $N$.
@@
    $\widetilde{\gamma}=(A^{-1}BA)\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)(A^{-1}BA)^{-1}=C\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)C^{-1}.$
@@
    and thus $(\widetilde{\gamma})^{p}$ is in $\widetilde{H}$. Hence either $\widetilde{h}=\widetilde{\gamma}$ if $\widetilde{\gamma}$ lies in $\mathrm{SL}_{2}(\mathbf{Z})$ or $\widetilde{h}=(\widetilde{\gamma})^{p}$. In particular, the cusp width at this cusp is either $n$ or $n/p$ and certainly divides $N$ and hence also $Np$.
@@ 4.2. Modular Forms
For an integer $N$, we will consider the following spaces of modular functions with rational coefficients generated by forms with bounded denominators, that is, subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}=\mathbf{Q}\llbracket q^{1/N}\rrbracket[1/q]$ (with $q=e^{\pi i\tau}$) generated by elements of $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes\mathbf{Q}$ as $\mathbf{Q}(\lambda)$-vector spaces.
@@
    Let $M_{2N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions on the modular curve $Y(2N)=\mathbf{H}/\langle E,\Gamma(2N)\rangle$ with coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$.
@@
    Let $R_{2N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions with coefficients in $\mathbf{Q}$, bounded denominators at the cusp $\zeta=i\infty$, and cusp widths dividing $2N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$.
(The vector space $R_{2N}$ was also defined in Proposition 3.0.1 but we repeat the definition here for convenience.) For example, the (weight 0) holomorphic modular forms on $Y(2)$ are given by $\mathbf{Q}[\lambda,1/\lambda,1/(1-\lambda)]$, and the $\mathbf{Q}(\lambda)$-vector space generated by such elements inside $\mathbf{Q}(\kern-1.49994pt{(}q)\kern-1.49994pt{)}$ with $q=e^{\pi i\tau}$ is $M_{2}=\mathbf{Q}(\lambda)$.
@@
There is a containment $M_{2N}\subset R_{2N}$, and $M_{2N}$ and $R_{2N}$ have finite dimensions over $M_{2}=\mathbf{Q}(\lambda)$.
@@
Let $f$ be a holomorphic modular function on $Y(2N)$, that is, a meromorphic function on the compact modular curve $X(2N)$ whose poles are all at the cusps. Assume also that $f$ has coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$, Then the modular form $f\Delta(\tau)^{m}$ is holomorphic at the cusps for sufficiently large $m$. Moreover, $f\Delta(\tau)^{m}$ has coefficients in $\mathbf{Q}$. It follows from [Shi71, Theorem 3.52] that $f\Delta(\tau)^{m}$ has bounded denominators. Since $\Delta^{-1}\in q^{-2}\mathbf{Z}\llbracket q\rrbracket$ has integral coefficients, it follows that $f$ also has bounded denominators, and thus there is a containment $M_{2N}\subset R_{2N}$.
The second claim follows from Corollary 2.0.5 and the remark (cf. the second paragraph of § 1.1.6) that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$. Indeed,
$\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$
is a well-defined holomorphic map with unit derivative at the origin, and hence by Schwarz’s lemma the universal covering $D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has derivative strictly larger than $1$ in absolute value. (Later, in Theorem 5.1.4 below, we will exactly compute this latter derivative.) ∎
@@
The vector spaces $M_{2N}$ and $R_{2N}$ are fields. The space $M_{2N}$ may be identified with the field of rational functions on the modular curve $Y(2N)/\mathbf{Q}$. There are injective algebra maps
$M_{2}\rightarrow M_{2N}\rightarrow R_{2N}.$
The space $R_{2N}$ is invariant under a normal finite index subgroup $G_{2N}\subset\langle E,\Gamma(2N)\rangle\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(G_{2N})=2N$.
@@
Note that $M_{2N}$ and $R_{2N}$ are subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$, which is a domain. Hence if $M_{2N}$ and $R_{2N}$ are rings then they are also integral domains, and any integral domain which has finite dimension over a field is itself a field.
The curve $Y(2N)$ has a standard model over $\mathbf{Q}$ (as a moduli space of elliptic curves $C$ with a given symplectic isomorphism $C[2N]\simeq\mathbf{Z}/2N\mathbf{Z}\oplus\mu_{2N}$) such that the cusp $\zeta=i\infty$ is defined over $\mathbf{Q}$, and the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$ on the global sections of $Y(2N)$ is compatible with the $q$-expansion map. It follows that the set of generators of $M_{2N}$ is closed under addition and multiplication and hence that $M_{2N}$ is a ring, and thus a field. Moreover, $M_{2N}$ contains the global sections of the (affine) curve $Y(2N)/\mathbf{Q}$, and hence $M_{2N}$ must be the function field of $Y(2N)/\mathbf{Q}$.
The vector space $R_{2N}$ is generated by holomorphic modular forms with bounded denominators at $\zeta=i\infty$. To show $R_{2N}$ is a ring, it suffices to show that the product of any two such generators $g$ and $h$ is also a generator. Certainly $gh$ is a holomorphic modular form with rational coefficients and bounded denominators, so it suffices to show that the cusp width still divides $2N$. But we may assume that $g$ and $h$ are invariant under finite index subgroup $G,H\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$, and thus $gh$ is invariant under $G\cap H$. It follows from Lemma 4.1.3 that $L(G\cap H)$ also has Wohlfahrt level dividing $2N$.
Since $R_{2N}$ is finite over $M_{2}$, it is generated by a finite number of basis elements each of which is invariant under some finite index subgroup $\Phi\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(\Phi)$ dividing $2N$. The intersection of all these groups still has finite index and level $2N$ by Lemma 4.1.3, and then we take $G_{2N}$ to be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in this intersection, which also has $L(G_{2N})=2N$ by Lemma 4.1.4. ∎